# DCNv3: Towards Next Generation Deep Cross Network for Click-Through Rate Prediction

### v6

@@ Abstract.
Deep & Cross Network and its derivative models have become an important paradigm for click-through rate (CTR) prediction due to their effective balance between computational cost and performance. However, these models face four major limitations: (1) the performance of existing explicit feature interaction methods is often weaker than that of implicit deep neural network (DNN), undermining their necessity; (2) many models fail to adaptively filter noise while increasing the order of feature interactions; (3) the fusion methods of most models cannot provide suitable supervision signals for their different sub-networks; (4) while most models claim to capture high-order feature interactions, they often do so implicitly and non-interpretably through DNN, which limits the trustworthiness of the model’s predictions.
To address the identified limitations, this paper proposes the next generation deep cross network: Deep Cross Network v3 (DCNv3), along with its two sub-networks: Linear Cross Network (LCN) and Exponential Cross Network (ECN) for CTR prediction. DCNv3 ensures interpretability in feature interaction modeling while linearly and exponentially increasing the order of feature interactions to achieve genuine Deep Crossing rather than just Deep & Cross. Additionally, we employ a Self-Mask operation to filter noise and reduce the number of parameters in the Cross Network by half. In the fusion layer, we use a simple yet effective multi-loss trade-off and calculation method, called Tri-BCE, to provide appropriate supervision signals. Comprehensive experiments on six datasets demonstrate the effectiveness, efficiency, and interpretability of DCNv3. The code, running logs, and detailed hyperparameter configurations are available at: https://github.com/salmon1802/DCNv3.
@@ 1. Introduction
Figure 1. Comparison among ECN, DCNv3, and other models in terms of network parameter number, AUC, and running time on Criteo dataset. $\star$ represents models using only explicit feature interactions, and $\circ$ represents models integrating implicit feature interactions. The graphic area represents the running time per epoch for each model (a larger area indicates a longer time, and vice versa).
Click-through rate (CTR) prediction is an essential part of industrial recommender systems (Zhu et al., 2021, 2022a). It uses user profiles, item attributes, and context to predict the probability of user-item interactions, thereby providing a better user experience and increasing the profitability of the recommender system (Wang et al., 2021b; Li et al., 2023; Chen et al., 2021; Guo et al., 2017). Most feature interaction-based CTR prediction models follow the paradigm proposed by DCN (Wang et al., 2017, 2021b), which aims to construct both explicit and implicit feature interactions and fuse the predictions of different interaction information to enhance interpretability and accuracy.
@@
  • Limited necessity for explicit interactions. As observed in Figure 1, most models (e.g., CIN (Lian et al., 2018)) using only explicit feature interactions achieve AUC performance below 81.3, while deep neural network (Covington et al., 2016) (DNN) achieves a better AUC of 81.4. This indicates that the performance of most explicit modeling methods is weaker than that of implicit DNN, which undoubtedly undermines the necessity of integrating explicit feature interactions. However, some works (Rendle et al., 2020; Wang et al., 2017) also highlight the limitations of DNN in capturing multiplicative feature interactions. Therefore, exploring a more effective method for explicit feature interactions is necessary.
  • Ineffective noise filtering capability. Many studies (Liu et al., 2020; Zhu et al., 2023a; Deng et al., 2021; Wang et al., 2023a) point out that CTR models contain a significant amount of redundant feature interactions and noise, especially in high-order feature interactions. Consequently, most CTR models (Song et al., 2019; Wang et al., 2021b; Lian et al., 2018) are built with only two or three network layers, abandoning the explicit capture of effective high-order feature interaction information. Meanwhile, filtering noise for the model often incurs additional computational costs, which can lead to longer training and inference times, potentially offsetting the benefits gained from improved model accuracy.
  • Insufficient and undifferentiated supervision signals. Most models using both explicit and implicit feature interaction methods require a fusion layer to obtain the final prediction (Wang et al., 2017, 2021b; Mao et al., 2023). However, they only use the final prediction to compute the loss, rather than providing appropriate supervision signals for the different methods themselves. This weakens the effectiveness of the supervision signals. Additionally, some studies (Wang et al., 2023b; Liu et al., 2024), such as the CL4CTR (Wang et al., 2023b) in Figure 1, attempt to introduce auxiliary loss to provide extra supervision signals. However, this often introduces additional computational costs and loss trade-off hyperparameters, increasing the difficulty of hyperparameter tuning. Therefore, a simple, efficient, and effective method for computing supervision signals is crucial.
@@
To address these limitations, this paper proposes the next generation deep cross network: Deep Cross Network v3 (DCNv3), along with its two sub-networks: Linear Cross Network (LCN) and Exponential Cross Network (ECN). DCNv3 integrates both low-order and high-order feature interactions while ensuring model interpretability by avoiding the use of DNN for implicit high-order feature interaction modeling. Specifically, we introduce a new exponentially growing Deep Crossing method to explicitly model high-order feature interactions, and categorize previous crossing methods as Shallow Crossing. Meanwhile, we design a Self-Mask operation to filter noise and reduce the number of parameters in the Cross Network by half. In the fusion layer, we propose a simple yet effective multi-loss calculation method, called Tri-BCE, to provide suitable supervision signals for different sub-networks.
@@
  • To the best of our knowledge, this is the first work to achieve surprising performance using only explicit feature interaction modeling without integrating DNN, which may contrast with the popular paradigms in the past CTR prediction literature. Therefore, this work may inspire further revisiting and innovation in feature interaction modeling.
  • We introduce a novel feature interaction modeling method, called Deep Crossing, which grows exponentially with the number of layers to achieve a genuine deep cross network. This method explicitly captures feature interaction information while using a Self-Mask operation to reduce the number of parameters by half and filter noise.
  • We propose a novel CTR model, called DCNv3, which explicitly captures both low-order and high-order feature interactions through two sub-networks, LCN and ECN. Additionally, we introduce a simple yet effective multi-loss trade-off and calculation method, called Tri-BCE, to ensure that different sub-networks receive appropriate supervision signals.
  • Comprehensive experiments on six datasets demonstrate the effectiveness, efficiency, and interpretability of DCNv3. Based on our experimental results, our models achieve 1st rankings on multiple CTR prediction benchmarks.
@@ 2.1. CTR Prediction
Explicit feature interactions are often modeled directly through hierarchical structures, such as the Cross Layer in the DCN (Wang et al., 2017), the Graph Layer in FiGNN (Li et al., 2019), and the Interacting Layer in AutoInt (Song et al., 2019). These methods ensure partial interpretability while allowing the capture of finite-order feature interactions. On the other hand, implicit feature interactions use DNN to automatically learn complex, non-manually defined patterns and interactions among features, further enhancing model performance (Wang et al., 2021a; Cheng et al., 2020; Wang et al., 2021b).
However, as the performance of explicit feature interactions is generally weaker than that of implicit feature interactions (Mao et al., 2023), several models attempt to abandon standalone explicit interaction methods and instead integrate multiplicative operations into DNN. MaskNet (Wang et al., 2021a) introduces multiplicative operations block by block, while GateNet (Huang et al., 2020), PEPNet (Chang et al., 2023), and FINAL (Zhu et al., 2023b) introduce them layer by layer to achieve higher performance. However, solely pursuing implicit modeling leads to a lack of interpretability. Meanwhile, most models lack the ability to filter noise and obtain appropriate supervisory signals. This paper aims to address these limitations through our proposed methods.
@@ 2.2. Problem Definition
A positive sample when $y=1$ and a negative sample when $y=0$. A CTR prediction model aims to predict $y$ and rank items based on the predicted probabilities $\hat{y}$.
@@
Implicit feature interaction aims to automatically learn complex non-manually defined data patterns and high-order feature interactions using DNN. It is highly efficient, performs well, but lacks interpretability.
Explicit feature interaction aims to model the combinations and relationships between input features directly through predefined functions, thereby improving model interpretability. A popular method for linearly increasing explicit feature interaction is $X_{n}=X_{1}\odot X_{n-1}$ (Wang et al., 2017, 2021b, 2023a). This method uses the Hadamard Product to interact feature $X_{n-1}$ with $X_{1}$ to generate the n-$th$ order feature $X_{n}$. Meanwhile, an exponentially growing feature interaction method is $X_{2^{n}}=X_{2^{n-1}}\odot X_{2^{n-1}}$, which can generate high-order features more efficiently.
@@ 3. Proposed Architecture
To address the limitations mentioned in Section 1, we propose the DCNv3 model, shown in Figure 2, which integrates different explicit feature interaction sub-networks: LCN and ECN, enabling it to simultaneously capture both low-order and high-order feature interactions while ensuring the model’s interpretability.
Figure 2. Architecture of the DCNv3.
@@ 3.1. Embedding & Reshape Layer
The input $X$ of the CTR prediction task, which is multi-field categorical data and is represented using one-hot encoding. Most CTR prediction models (Song et al., 2019; Li et al., 2023; Sang et al., 2024) utilize an embedding layer to transform them into low-dimensional dense vectors: $\mathbf{e}_{i}=\textit{E}_{i}x_{i}$, where $\textit{E}_{i}\in\mathbb{R}^{d\times s_{i}}$ and $s_{i}$ separately indicate the embedding matrix and the vocabulary size for the $i$-th field, $d$ represents the embedding dimension. In our model, to enable the network to obtain richer information from multi-views, we reshape the embeddings using the chunk operation (Li et al., 2024a) to divide them into two different views: $\mathbf{e}_{i,a},\mathbf{e}_{i,b}=\texttt{chunk}(\mathbf{e}_{i})$, where $\mathbf{e}_{i,a}$ is considered an original view, and $\mathbf{e}_{i,b}$ is considered another view. Further, we can get the resulting embedding after reshaping: $\mathbf{x_{1}}=\left[\mathbf{e}_{1,a},\cdots,\mathbf{e}_{f,a},\mathbf{e}_{1,b},\cdots,\mathbf{e}_{f,b}\right]\in\mathbb{R}^{D}$, where $f$ denotes the number of fields, $D=\sum_{i=1}^{f}d$, and $\mathbf{x_{1}}$ denotes the first-order features shared as input by LCN and ECN.
3.2. Deep Cross Network v3
LCN: It is used for low-order (shallow) explicit feature interaction with linear growth. Its recursive formula is as follows:
(2) $\mathbf{c}_{l}$ $=\mathbf{W}_{l}\mathbf{x}_{l}+\mathbf{b}_{l},$
@@ 3.2. Deep Cross Network v3
ECN: As the core structure of DCNv3, it is used for high-order (deep) explicit feature interaction with exponential growth. Its recursive formula is as follows:
(3) $\mathbf{c}_{l}$ $=\mathbf{W}_{l}\mathbf{x}_{2^{l-1}}+\mathbf{b}_{l},$
@@
Figure 3. Visualization of different crossing methods. $\times$ denotes matrix multiplication. The left side shows the workflow of Cross Network v2 (Wang et al., 2021b), while the right side shows the workflow of DCNv3. To simplify the representation and facilitate understanding of their differences, we use the same math notation.
Self-Mask: Figure 3 visualizes the computation process of different crossing methods. The Cross Vector $\mathbf{c}_{l}$ aims to model interactions between features at the bit-level, while the weight matrix $\mathbf{W}_{l}$ aims to compute the inherent importance of different feature fields. However, as pointed out by some works (Xiao et al., 2017; Liu et al., 2020; Wang et al., 2023a), not all feature interactions are beneficial for the final prediction in CTR tasks. Therefore, we introduce the Self-Mask operation to filter out noisy information from feature interactions in another view (corresponding to $\texttt{Mask}(\mathbf{c}_{l})$) while maintaining the integrity of the interaction information of the original view (corresponding to $\mathbf{c}_{l}$). The Self-Mask operation is as follows:
@@
Fusion: Most previous CTR models (Zhu et al., 2023b; Mao et al., 2023; Wang et al., 2021b) attempt to model explicit and implicit feature interactions, which essentially means capturing both low-order and high-order feature interactions. Our DCNv3 achieves this by integrating LCN and ECN, avoiding the use of less interpretable DNN:
(5) $\hat{y}_{\scriptscriptstyle D}$ $=\sigma(\mathbf{W}_{\scriptscriptstyle D}\mathbf{x}_{{2^{L}}}+\mathbf{b}_{\scriptscriptstyle D}),$
$\hat{y}_{\scriptscriptstyle S}$ $=\sigma(\mathbf{W}_{\scriptscriptstyle S}\mathbf{x}_{L+1}+\mathbf{b}_{\scriptscriptstyle S}),$
$\hat{y}$ $=\texttt{Mean}(\hat{y}_{\scriptscriptstyle D},\hat{y}_{\scriptscriptstyle S}),$
where $\mathbf{W}_{\scriptscriptstyle D}$ and $\mathbf{W}_{\scriptscriptstyle S}\in\mathbb{R}^{1\times D}$ represent learnable weights, $\mathbf{b}_{\scriptscriptstyle D}$ and $\mathbf{b}_{\scriptscriptstyle S}$ are biases, Mean denotes the mean operation, $\hat{y}_{\scriptscriptstyle D},\hat{y}_{\scriptscriptstyle S}$, $\hat{y}$ represent the prediction results of ECN, LCN, and DCNv3, respectively, and $L$ denotes the last number of layers.

### v7

# FCN: Fusing Exponential and Linear Cross Network for Click-Through Rate Prediction

@@ Abstract.
As an important modeling paradigm in click-through rate (CTR) prediction, the Deep & Cross Network (DCN) and its derivative models have gained widespread recognition primarily due to their success in a trade-off between computational cost and performance. This paradigm employs a cross network to explicitly model feature interactions with linear growth, while leveraging deep neural networks (DNN) to implicitly capture higher-order feature interactions. However, these models still face several key limitations: (1) The performance of existing explicit feature interaction methods lags behind that of implicit DNN, resulting in overall model performance being dominated by the DNN; (2) While these models claim to capture high-order feature interactions, they often overlook potential noise within these interactions; (3) The learning process for different interaction network branches lacks appropriate supervision signals; and (4) The high-order feature interactions captured by these models are often implicit and non-interpretable due to their reliance on DNN.
To address the identified limitations, this paper proposes a novel model, called Fusing Cross Network (FCN), along with two sub-networks: Linear Cross Network (LCN) and Exponential Cross Network (ECN). FCN explicitly captures feature interactions with both linear and exponential growth, eliminating the need to rely on implicit DNN. Moreover, we introduce the Self-Mask operation to filter noise layer by layer and reduce the number of parameters in the cross network by half. To effectively train these two cross networks, we propose a simple yet effective loss function called Tri-BCE, which provides tailored supervision signals for each network. We evaluate the effectiveness, efficiency, and interpretability of FCN on six benchmark datasets. Furthermore, by integrating LCN and ECN, FCN achieves a new state-of-the-art performance. The code, running logs, and detailed hyperparameter configurations are available at: https://github.com/salmon1802/FCN/tree/master.
@@ 1. Introduction
Click-through rate (CTR) prediction is an essential part of industrial recommender systems (Zhu et al., 2021, 2022a). It uses user profiles, item attributes, and context to predict the probability of user-item interactions, thereby providing a better user experience and increasing the profitability of the recommender system (Wang et al., 2021b; Li et al., 2023; Chen et al., 2021; Guo et al., 2017). Most feature interaction-based CTR prediction models follow the paradigm proposed by DCN (Wang et al., 2017, 2021b), which aims to construct both explicit low-order and implicit high-order feature interactions and fuse the predictions of different interaction information to enhance interpretability and accuracy (Lian et al., 2018; Wang et al., 2021b).
Figure 1. Comparison among ECN, FCN, and other models in terms of network parameter number, AUC, and running time on Criteo dataset. $\star$ represents models using only explicit feature interactions, and $\circ$ represents models integrating implicit feature interactions. The graphic area represents the running time per epoch for each model (a larger area indicates a longer time, and vice versa).
@@
  • Limited necessity for explicit interactions. As observed in Figure 1, most models (e.g., CIN (Lian et al., 2018)) using only explicit feature interactions achieve AUC performance below 81.3, while deep neural network (Covington et al., 2016) (DNN) achieves a better AUC of 81.4. This indicates that the performance of most explicit modeling methods is weaker than that of implicit DNN. It raises an intriguing question: if implicit feature interaction performs so well, why not abandon explicit interaction methods altogether and rely solely on implicit modeling with DNNs? FinalMLP (Mao et al., 2023), also shown in Figure 1, answers this question by demonstrating strong performance with a two-stream MLP architecture. This result further challenges the necessity of integrating explicit feature interactions in such models. However, some work (Rendle et al., 2020; Wang et al., 2017) also highlights the limitations of DNN in capturing multiplicative feature interactions, particularly in polynomial fitting (Wang et al., 2021b). Therefore, exploring a more effective method for explicit feature interactions is necessary.
  • Ineffective noise filtering capability. Many studies (Liu et al., 2020; Zhu et al., 2023a; Deng et al., 2021; Wang et al., 2023a) point out that CTR models contain a significant amount of redundant feature interactions and noise, especially in high-order feature interactions. Consequently, most CTR models (Song et al., 2019; Wang et al., 2021b; Lian et al., 2018) are built with only two or three network layers, abandoning the explicit capture of effective high-order feature interaction information. Meanwhile, filtering noise for the model often incurs additional computational costs, which can lead to longer training and inference times (Wang et al., 2023a), potentially offsetting the benefits gained from improved model accuracy.
  • Insufficient and undifferentiated supervision signals. Most models using both explicit and implicit feature interaction methods require a fusion layer to obtain the final prediction (Wang et al., 2017, 2021b; Mao et al., 2023). However, they only use the final prediction to compute the loss, rather than providing appropriate supervision signals for the different methods themselves. This weakens the effectiveness of the supervision signals. Additionally, some studies (Wang et al., 2023b; Liu et al., 2024), such as the CL4CTR (Wang et al., 2023b) in Figure 1, attempt to introduce auxiliary loss to provide extra supervision signals. Nonetheless, this often introduces additional computational costs and loss balance hyperparameters, increasing the difficulty of hyperparameter tuning. Therefore, a simple, efficient, and effective method for computing supervision signals is crucial.
@@
To address these limitations, this paper proposes a novel model, called Fusing Cross Network (FCN) for CTR prediction. FCN enhances high-order feature interactions and their interpretability through two sub-networks: Linear Cross Network (LCN) and Exponential Cross Network (ECN). LCN and ECN progressively increase the order of feature interactions linearly and exponentially, respectively. Specifically, we propose a new exponentially growing Deep Crossing method to explicitly model high-order feature interactions, and categorize previous crossing methods (Wang et al., 2021b) as Shallow Crossing. Meanwhile, we design a Self-Mask operation to filter noise and reduce the number of parameters in the Cross Network by half. In the fusion layer, we propose a simple yet effective loss function, called Tri-BCE, to provide suitable supervision signals for different sub-networks.
@@
  • To the best of our knowledge, this is the first work to achieve surprising performance using only explicit feature interaction modeling without integrating DNN, which may contrast with the popular paradigms in the past CTR prediction literature.
  • We propose a novel explicit high-order feature interaction method that captures exponentially growing feature interaction information, effectively filters noise through a self-masking operation, and reduces the number of parameters by half.
  • We propose a novel CTR model, called FCN, which explicitly captures both low-order and high-order feature interactions through two sub-networks, LCN and ECN. Additionally, we introduce a simple yet effective multi-loss balance and calculation method, called Tri-BCE, to ensure that different sub-networks receive appropriate supervision signals.
  • Comprehensive experiments on six datasets demonstrate the effectiveness, efficiency, and interpretability of FCN. Based on our experimental results, our models achieve 1st rankings on multiple CTR prediction benchmarks.
@@ 2.1. CTR Prediction
Explicit feature interactions are often modeled directly through hierarchical structures, such as the Cross Layer in the DCN (Wang et al., 2017), the Graph Layer in FiGNN (Li et al., 2019), and the Interacting Layer in AutoInt (Song et al., 2019). These methods ensure partial interpretability while allowing the capture of finite-order feature interactions. On the other hand, some studies attempt to integrate implicit feature interactions by designing different structures. These structures mainly include stacked structures (Yu et al., 2021; Qu et al., 2016, 2018), parallel structures (Zhu et al., 2023b; Wang et al., 2023a; Li et al., 2023), and alternate structures (Li et al., 2024a; Zhu et al., 2023b). The introduction of these structures not only enhances the expressive power of the models but also captures high-order feature interactions through DNN, leading to significant performance improvements in practical applications.
However, as the performance of explicit feature interactions is generally weaker than that of implicit feature interactions (Mao et al., 2023), several models attempt to abandon standalone explicit interaction methods and instead integrate multiplicative operations into DNN. MaskNet (Wang et al., 2021a) introduces multiplicative operations block by block, while GateNet (Huang et al., 2020), PEPNet (Chang et al., 2023), and FINAL (Zhu et al., 2023b) introduce them layer by layer to achieve higher performance. Nevertheless, the use of implicit feature interactions leads to a lack of interpretability. Meanwhile, most models lack the ability to filter noise and obtain appropriate supervision signals. This paper aims to address these limitations through our proposed methods.
@@ 2.2. Problem Definition
where $y=1$ represents a positive sample and $y=0$ represents a negative sample. A CTR prediction model aims to predict $y$ and rank items based on the predicted probabilities $\hat{y}$.
@@
Implicit feature interaction aims to automatically learn complex non-manually defined data patterns and high-order feature interactions using DNN (Lian et al., 2018; Wang et al., 2021b). Formally, for a given input feature $X^{1}$, the implicit feature interaction process can be expressed as:
(2) $X^{k+1}=\sigma\left(\mathbf{W}^{k}X^{k}+\mathbf{b}^{k}\right),\ \ k=1,2,\dots,K,$
where $k$ denotes the depth of the layer and $\sigma$ represents the activation function. Compared to explicit feature interaction, implicit feature interaction does not have a concrete interaction form. However, it is highly efficient and performs well, but lacks interpretability.
Explicit feature interaction aims to model the combinations and relationships between input features directly through predefined functions with a controllable order, thereby improving model interpretability. A popular method for linearly increasing explicit feature interaction is $X_{n}=X_{1}\odot X_{n-1}$ (Wang et al., 2017, 2021b, 2023a). This method uses the Hadamard Product to interact feature $X_{n-1}$ with $X_{1}$ to generate the n-$th$ order feature $X_{n}$. In this paper, we propose an exponentially growing feature interaction method: $X_{2^{n}}=X_{2^{n-1}}\odot X_{2^{n-1}}$, which can generate high-order features more efficiently.
@@ 3. Proposed Architecture
To address the limitations mentioned in Section 1, we propose the FCN model, shown in Figure 2, which integrates different explicit feature interaction sub-networks: LCN and ECN, enabling it to simultaneously explicitly capture both low-order and high-order feature interactions while ensuring the model’s interpretability.
Figure 2. Architecture of the FCN.
@@ 3.1. Embedding & Reshape Layer
The input $X$ of the CTR prediction task, which is multi-field categorical data and is represented using one-hot encoding. Most CTR prediction models (Song et al., 2019; Li et al., 2023; Sang et al., 2024) utilize an embedding layer to transform them into low-dimensional dense vectors: $\mathbf{e}_{i}=\textit{E}_{i}x_{i}$, where $\textit{E}_{i}\in\mathbb{R}^{d\times s_{i}}$ and $s_{i}$ separately indicate the embedding matrix and the vocabulary size for the $i$-th field, $d$ represents the embedding dimension. In our model, to enable the network to obtain richer information from multi-views, we reshape the embeddings using the chunk operation (Li et al., 2024a) to divide them into two different views: $\mathbf{e}_{i,a},\mathbf{e}_{i,b}=\texttt{chunk}(\mathbf{e}_{i})$, where $\mathbf{e}_{i,a}$ is considered an original view, and $\mathbf{e}_{i,b}$ is considered another view. Further, we can get the resulting embedding after reshaping: $\mathbf{x_{1}}=\left[\mathbf{e}_{1,a},\cdots,\mathbf{e}_{f,a},\mathbf{e}_{1,b},\cdots,\mathbf{e}_{f,b}\right]\in\mathbb{R}^{D}$, where $f$ denotes the number of fields, $D=\sum_{i=1}^{f}d$, and $\mathbf{x_{1}}$ denotes the first-order features [^1] In this work, to ensure that the number of network layers aligns with the order of feature interactions, we take the input $\mathbf{x_{1}}$ as the first-order feature interactions shared as input by LCN and ECN.
3.2. Fusing Cross Network
Previous deep CTR models (Wang et al., 2021b; Lian et al., 2018; Guo et al., 2017; Wang et al., 2023a) typically use explicit interaction methods to capture low-order feature interactions and rely on DNN to implicitly capture high-order feature interactions. However, the former is limited to capturing low-order relationships due to complexity constraints and generally exhibits lower performance (Mao et al., 2023), while the latter depends on the nonlinear learning of deep networks, making it difficult to provide clear interpretability of feature interactions. Therefore, to ensure overall model interpretability while effectively capturing high-order feature interactions, we attempt to design an explicit feature interaction modeling method with exponential growth, eliminating the need to introduce DNN, thereby achieving both performance and interpretability.
3.2.1. Linear Cross Network (LCN)
It is used for low-order (shallow) explicit feature interaction with linear growth. Its recursive formula is as follows:
(3) $\mathbf{c}_{l}$ $=\mathbf{W}_{l}\mathbf{x}_{l}+\mathbf{b}_{l},\ \ l=1,2,\dots,L,$
@@ 3.2.1. Linear Cross Network (LCN)
3.2.2. Exponential Cross Network (ECN)
As the core structure of FCN, it is used for high-order (deep) explicit feature interaction with exponential growth. Its recursive formula is as follows:
(4) $\mathbf{c}_{l}$ $=\mathbf{W}_{l}\mathbf{x}_{2^{l-1}}+\mathbf{b}_{l},\ \ l=1,2,\dots,L,$
@@ 3.2.2. Exponential Cross Network (ECN)
Figure 3. Visualization of different crossing methods. $\times$ denotes matrix multiplication. The left side shows the workflow of Cross Network v2 (Wang et al., 2021b), while the right side shows the workflow of FCN. To simplify the representation and facilitate understanding of their differences, we use the same math notation.
3.2.3. Self-Mask
Figure 3 visualizes the computation process of different crossing methods. The Cross Vector $\mathbf{c}_{l}$ aims to model interactions between features at the bit-level, while the weight matrix $\mathbf{W}_{l}$ aims to compute the inherent importance of different feature fields. However, as pointed out by some works (Xiao et al., 2017; Liu et al., 2020; Wang et al., 2023a), not all feature interactions are beneficial for the final prediction in CTR tasks. Therefore, we introduce the Self-Mask operation to filter out noisy information from feature interactions in another view (corresponding to $\texttt{Mask}(\mathbf{c}_{l})$) while maintaining the integrity of the interaction information of the original view (corresponding to $\mathbf{c}_{l}$). The Self-Mask operation is as follows:
@@ 3.2.3. Self-Mask
3.2.4. Fusion Sub-networks
Most previous CTR models (Zhu et al., 2023b; Mao et al., 2023; Wang et al., 2021b) attempt to model explicit and implicit feature interactions, which essentially means capturing both low-order and high-order feature interactions. Our FCN achieves this by integrating LCN and ECN, avoiding the use of less interpretable DNN:
(6) $\hat{y}_{\scriptscriptstyle\mathbb{D}}$ $=\sigma(\mathbf{W}_{\scriptscriptstyle\mathbb{D}}\mathbf{x}_{{2^{L}}}+\mathbf{b}_{\scriptscriptstyle\mathbb{D}}),$
$\hat{y}_{\scriptscriptstyle\mathbb{S}}$ $=\sigma(\mathbf{W}_{\scriptscriptstyle\mathbb{S}}\mathbf{x}_{L+1}+\mathbf{b}_{\scriptscriptstyle\mathbb{S}}),$
$\hat{y}$ $=\texttt{Mean}(\hat{y}_{\scriptscriptstyle\mathbb{D}},\hat{y}_{\scriptscriptstyle\mathbb{S}}),$
where $\mathbf{W}_{\scriptscriptstyle\mathbb{D}}$ and $\mathbf{W}_{\scriptscriptstyle\mathbb{S}}\in\mathbb{R}^{1\times D}$ represent learnable weights, $\mathbf{b}_{\scriptscriptstyle\mathbb{D}}$ and $\mathbf{b}_{\scriptscriptstyle\mathbb{S}}$ are biases, Mean denotes the mean operation, $\hat{y}_{\scriptscriptstyle\mathbb{D}},\hat{y}_{\scriptscriptstyle\mathbb{S}}$, $\hat{y}$ represent the prediction results of ECN, LCN, and FCN, respectively, and $L$ denotes the last number of layers.